OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -4x - y < -6 A, Dashed line with shading on the side that includes the origin. B, Solid line with shading on the side that does not include the origin. C, Dashed line with shading on the side that does not include the origin. D, Solid line with shading on the side that includes the origin.

OpenStudy (compassionate):

y > -4x + 6 Dashed line. 0 > -4(0) + 6 0 > 6 The side does not include the origin. As always, medals are appreciated.

OpenStudy (anonymous):

im not under standing

OpenStudy (compassionate):

What part don't you understand, hun?

OpenStudy (anonymous):

all im sorry this is my first time doing any of this

OpenStudy (compassionate):

Okay, well, let me know what parts you need help with. Are you confused with the setting up the equation? Do you know what y = mx + b is?

OpenStudy (anonymous):

setting up the equation and y = mx + b is

OpenStudy (anonymous):

what im not under standing

OpenStudy (compassionate):

Okay, so the current way the equation is set up is called Standard Form. Standard form is not good for graphing because it is confusing. y = mx + b is an easy way to graph equations. When someone is in y = mx + b, m = slope and b = y-intercept. For example, if I had y = 2x + 1, I would go to 1 on my y-coordinate and up two, and over one. Check this out: http://www.purplemath.com/modules/strtlneq.htm